OpenStudy (anonymous):

y=(a-ib-a)2+b2 y is supposed to equal 0. how? I get it to 2b^2 i^2=-1

OpenStudy (anonymous):

confused is this \[y=(a+ib-a)^2b^2\]?

OpenStudy (anonymous):

\[y=(a+ib-a)^2+b^2\]

OpenStudy (anonymous):

f(x)=(x-a)^2+b^2 g(x)=-(x-a)^2+b^2

OpenStudy (anonymous):

then you get 0 for sure because \[a-a=0\] and you have \[y=(bi)^2+b^2=i^2b^2+b^2=-b^2+b^2=0\]

OpenStudy (anonymous):

yes but f(x) has zeros at a+-b

OpenStudy (anonymous):

\[f(x)=(x-a)^2+b^2\] \[ g(x)=-(x-a)^2+b^2\]what is the question?

OpenStudy (anonymous):

prove that f(x) has its zeros at a±ib where i=√(-1).

OpenStudy (anonymous):

ah ok then write \[(x-a)^2=-b^2\] as a first step

OpenStudy (anonymous):

then \[x-a=\pm\sqrt{-b^2}=\pm\sqrt{-1}\sqrt{b^2}=\pm bi\]

OpenStudy (anonymous):

and so \[x=a\pm bi\]

OpenStudy (anonymous):

aha makes sense

OpenStudy (anonymous):

but then, how can i prove that f(x)+g(x)=0?

OpenStudy (anonymous):

hmm i don't think it is zero

OpenStudy (anonymous):

its supposed to be :(

OpenStudy (anonymous):

\[f(x)+g(x)=(x-a)^2+b^2+-(x-a)^2+b^2=2b^2\]

OpenStudy (anonymous):

possibly theres a complez zero at the height 2b^2

OpenStudy (anonymous):

\[f(x)+g(x)=2b^2\] is a constant. not zero unless b is zero

OpenStudy (anonymous):

no, there are no complex zeros. \[f(x)+g(x)\] is a constant, there is no x in it

OpenStudy (anonymous):

what does the question say exactly?

OpenStudy (anonymous):

well we have a formula for quartics and one for cubics and i need to find a general statement true for both.

OpenStudy (anonymous):

??

OpenStudy (anonymous):

Cubics: y1=(x+2)(x-(3+2i))(x+(3-2i))=x3-4x2+x+26 y2=-(x+2)(x-(3+2))(x-(3-2))=-x3+4x2+7x-10 ym=4x+8 y1+y2=8x+16=2(4x+8)=2ym

OpenStudy (anonymous):

my statement is that f(x)+g(x)=2ym

OpenStudy (anonymous):

what does ym mean?

OpenStudy (anonymous):

thats a straight line which has an equal vertical distance to f(x) and g(x) at all points.

OpenStudy (anonymous):

in the quartics case it is the line y=0

OpenStudy (anonymous):

ok so in simple language y = f(x) + g(x) ?

OpenStudy (anonymous):

well 2y=f(x)+g(x)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so now you have \[f(x)=(x-a)^2+b^2,g(x)=-(x-a)^2+b^2\] and \[f(x)+g(x)=2b^2\] right?

OpenStudy (anonymous):

btw these are quadratics not "quartics" because the degree is 2, not 4

OpenStudy (anonymous):

ah i got confused, the straight line ym isnt at the height 0, its at b^2

OpenStudy (anonymous):

ooooooooooooh i think i see the line that is equidistant from these two is the line y = 0

OpenStudy (anonymous):

i dont know what equidistant means unfortunately haha

OpenStudy (anonymous):

omg hold on this is such a stupid question

OpenStudy (anonymous):

of course f(x)+g(x)=2b^2... b^2=ym

OpenStudy (anonymous):

they share a vertex

OpenStudy (anonymous):

right they sure do

OpenStudy (anonymous):

i get it now though haha thanks for your time ^^

OpenStudy (anonymous):

and that is (a,b^2) so they are equidistant from the line \[y=b^2\]

OpenStudy (anonymous):

yw